OpenStudy (babyslapmafro):

Can someone please help me find the derivative of the following function?

OpenStudy (babyslapmafro):

\[ y=\cos^7(\sin10x)\]

OpenStudy (babyslapmafro):

How do I go about finding the derivative?

OpenStudy (babyslapmafro):

\[ y'=\cos^7(\frac{ d}{ dx }\sin[10x])+(\sin[10x])(\frac{ d }{ dx }\cos^7)\]

OpenStudy (babyslapmafro):

Am I on the right track?

OpenStudy (cwrw238):

the 7 is an exponent right?

OpenStudy (cwrw238):

if so you use the chain rule

OpenStudy (babyslapmafro):

avoid the product rule altogether?

OpenStudy (cwrw238):

i think so

OpenStudy (cwrw238):

i'm not thinking straight today - i'll have another look at this

OpenStudy (babyslapmafro):

no worries, thanks

OpenStudy (cwrw238):

ok let u = sin 10x du/dx = 10 cos10x y = cos^7 u dy/du = -7 sin^6 u = -7 (sin sin10x)^6 dy/dx = -70 cos 10x (sin sin 10x)^6

OpenStudy (allank):

I would first write it out like (cos(sin10x))^7 to clearly see the 'outside' and 'inside' functions for chain rule application. So, stepwise: d/dx of (cos(sin 10x))^7 =7(cos(sin 10x))^6 * d/dx of cos(sin 10x) = .......................... * -sin(sin 10x) * d/dx of sin (10x) = .......................... * .................... * cos(10x) * d/dx of (10x) = .......................... * .................... * ............. * 10 Then putting that all together gives 7(cos(sin 10x))^6 * -sin(sin 10x) * cos(10x) * 10 You may do some simplification if you want to. I hope that helps.